OpenStudy (anonymous):

What is the equation of the line that passes through (0, -2) and has a slope of 0? y = -2 y = 2 x = 2

OpenStudy (tkhunny):

Slope of 0 (zero) is Horizontal. This rules out "x = " What else?

OpenStudy (anonymous):

so it would be y=-2 ?

OpenStudy (tkhunny):

Does it pass through the required point?

OpenStudy (anonymous):

(Y-(-2))=0*(x-0| y+2=0 y=-2

OpenStudy (tkhunny):

Don't try that with a Vertical Line.

OpenStudy (anonymous):

@bigeyes420 : Your answer is right !:)

OpenStudy (anonymous):

okay thank you :)